OpenStudy (sarahc):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)